### v1

@@ Abstract
Object detection on visible (RGB) and infrared (IR) images, as an emerging solution to facilitate robust detection for around-the-clock applications, has received extensive attention in recent years. With the help of IR images, object detectors have been more reliable and robust in practical applications by using RGB-IR combined information. However, existing methods still suffer from modality miscalibration and fusion imprecision problems. Since transformer has the powerful capability to model the pairwise correlations between different features, in this paper, we propose a novel Calibrated and Complementary Transformer called $\textbf{C}^{2}\textbf{Former}$ to address these two problems simultaneously. In $\textbf{C}^{2}\textbf{Former}$, we design an Inter-modality Cross-Attention (ICA) module to obtain the calibrated and complementary features by learning the cross-attention relationship between the RGB and IR modality. To reduce the computational cost caused by computing the global attention in ICA, an Adaptive Feature Sampling (AFS) module is introduced to decrease the dimension of feature maps. Because $\textbf{C}^{2}\textbf{Former}$ performs in the feature domain, it can be embedded into existed RGB-IR object detectors via the backbone network. Thus, one single-stage and one two-stage object detector both incorporating our $\textbf{C}^{2}\textbf{Former}$ are constructed to evaluate its effectiveness and versatility. With extensive experiments on the DroneVehicle and KAIST RGB-IR datasets, we verify that our method can fully utilize the RGB-IR complementary information and achieve robust detection results. The code is available at https://github.com/yuanmaoxun/Calibrated-and-Complementary-Transformer-for-RGB-Infrared-Object-Detection.git.
@@
RGB-infrared object detection, modality calibration, complementary fusion, remote sensing images, multispectral object detection.
@@ I Introduction
Object detection, as a core technology in the computer vision field, has been used in various practical applications, such as autonomous driving, video surveillance and robotics, etc. With the development of convolutional neural network (CNN), the past few years have witnessed significant progress in object detectors [1, 2, 3, 4, 5]. Since these detectors are mainly designed for visible (RGB) images, they cannot cope with the challenges in bad illumination (nighttime) or diverse weather conditions (rain, dust and fog) [6, 7]. For these reasons, detectors exploiting infrared images have been widely adopted as an additional modality for robust object detection [8, 9, 10]. To achieve the better performance in the full-time detection task, a growing number of works [11, 12, 13, 14, 15] have been investigated to obtain rich features of objects by integrating visible (RGB) cues and infrared (IR) spectra.
@@
However, there are still two major problems in RGB-IR object detection [17, 16]. (1) One is modality miscalibration as shown in Figure 1(a). Existing methods [18, 19, 13] usually assume that RGB-IR image pairs are perfectly geometrically aligned, and they directly perform the multi-modal fusion methods. Actually, after the image registration algorithm, the paired images are just weakly aligned [20, 21]. This because the RGB-IR image pairs are captured by different sensors with different field-of-views (FoVs) and imaging time, the positions of moving objects may be out of synchronization. (2) Another problem is fusion imprecision as shown in Figure 1(b). In IR modality, the thermal radiation of the vehicle can be important hints for the detector, but none of the RGB images has such cues in bad illumination. The RGB and IR modality features are diverse in terms of object color, texture and properties as they are captured in different spectral bands, which brings a big challenge to the fusion strategy. In fact, these two problems are tightly coupled, the misaligned features will directly lead to the imprecision fusion. Thus, these two problems need to be bound together and solved simultaneously.
@@
Recently, some approaches have been introduced to solve these problems in RGB-T detection task. Zhang et al. [20] and Yuan et al. [21] solve the modality miscalibration problem by predicting the proposals’ offset between two modalities in the RoI head. Although these instance-level prediction solves the object misalignment problem to some extent, they cannot guarantee that the features inside two modality proposals are perfectly aligned. Besides, some cross-modality fusion strategy [7, 22, 12] are also conducted to address the fusion imprecision problem through the simple feature concatenation or the illumination-aware fusion. However, the inherent complementary information between RGB-IR modality features is still not exploited. More importantly, the above methods do not consider correlating the problems of modality miscalibration and fusion imprecision, but mainly focus on one of them or solve them separately, which will affect the detection performance.
@@
In this paper, refer to the self-attention mechanism in the transformer [23], we formulate the aforementioned two problems into transformer framework and propose a novel Calibrated and Complementary TransFormer-based fusion method called $\textup{C}^{2}$ormer to these two problems simultaneously. Specifically, for each feature point in one modality (Query vector), we first compute the similarity score with all the feature points in the other modality (Key vectors), and then obtain the Attention values via the weighted sum versus the features in the other modality (Value vectors). These attention values have two properties: (1) they incorporate the global context information from the other modality, therefore, can be regarded as a better complementary feature than the single local feature point to the original modality. Thus, we can simply concatenate these attention values with the original feature point to conduct an effective feature fusion. (2) The most relevant feature points in the other modality contribute the most to the attention values, which implies that the calibration is automatically achieved thanks to the similarity score. In this way, we simultaneously perform the calibration and fusion operation in an unified framework. Based on this idea, we design an Inter-modality Cross-Attention (ICA) module. An illustration result of the ICA module is given in Figure 2.
@@
Directly computing the attention across global feature maps has high computational costs. For that, it needs a feature sampling to reduce the dimension of feature maps. However, because of the miscalibration, an uniform sampling strategy for both modalities may omit meaningful features. To tackle this issue, we present an Adaptive Feature Sampling (AFS) strategy. In detail, we predict an offset versus the foreground object across modalities through a deviation network, and then generate the corresponding sampling strategy according to this offset for different modalities. In this way, we can ensure the integrity for important regions in the sampled feature maps. In actual usage, the given feature maps are first down-sampled by AFS module, and then processed by ICA module to obtain the complementary features, which are finally concatenated with the given feature maps to obtain the fusion features.
@@
The rest of the article is organized as follows. In the next section we review the scientific literature related to our proposed approach. Section III analyses the problem and describes our approach. Experimental results are presented in Section IV, along with the comparison with state-of-the-art methods, and in Section V we conclude with a discussion of our contribution.
@@ II-A Multi-modal Transformers
Since transformer model [23] has shown its powerful ability to capture long-range dependencies, multi-modal transformers [24, 25, 26, 27] have also been explored. To fuse different modality features, CMSA [28] feeds the words and images into the transformer to construct a joint multimodal feature representation for referring image segmentation and BrT [26] uses a transformer to allow interactive learning by adding point-to-patch projection between images and point clouds. Similarly, Liu et al. [24] and Xiao et al. [27] design a triplet transformer and feature exchange transformer respectively to detect the salient objects in RGB-D images. However, these methods perform feature fusion by directly concatenating or summing the multi-modalities to feed the transformer. They do not consider feature misalignment problem, which will lead to limited representation ability of the fused features.
@@
Different from above approaches, Pixel-BERT [29] considers the modality misalignment problem and align vision and language semantics in pixel and text level using self-supervised learning. Besides, Wang et al. [30] leverages the hierarchical alignment to explore the detailed and diverse characteristics in text-video retrieval task. Recently, Wang et al. [31] aggregate RGB and depth cues through projecting alignment features from other modalities at the token level. Even though these methods consider the feature misalignment problem, but the modality complementarity is not further exploited in the fusion process.
@@
Above approaches attempt to utilize the transformer to address multi-modal problems. However, compared with these multi-modal tasks (words-images, texts-videos and point clouds-images), our $\textup{C}^{2}$ormer is designed specific to the RGB-T object detection task and we are the first apply transformer to this task. Since the problems of modality miscalibration and fusion imprecision are unique and intractable in RGB-T image pairs, we are inspired by the core idea of the self-attention mechanism and formulate these two problems into transformer to establish their correlation. Thus, these two problems are combined together and solved simultaneously, which is of great significance in RGB-T object detection task.
II-B Oriented Object Detection
@@
Oriented object detection refers to using rotated bounding box to represent building detectors, which can provide more accurate location of objects. Xia et al. [32] construct a large-scale object detection benchmark with oriented annotations, named DOTA. Existing methods of building detectors based on horizontal bounding boxes are plagued by multiple objects of interest in one anchor. Therefore, many researches about oriented object detection [33, 34, 35, 36, 37] have been presented to alleviate this problem. Ma et al. [33] construct R-RPN to detect oriented objects by rotated proposals. Jiang et al. present $\rm{R^{2}}$ [34], which has a horizontal region of interest (RoI) that is leveraged to predict both horizontal and rotated boxes. To avoid a large number of anchors, RoI Transformer [35] is proposed to transform the horizontal RoI in $\rm{R^{2}}$ into a rotated one. To speed up the inference time of detectors, some works [38, 39, 40, 41] explore one-stage and anchor-free oriented object detectors. $\rm R^{3}$et [38] and $\rm S^{2}$-Net [39] adopt alignments between horizontal receptive fields and rotated anchors to get a better feature representation. Recently, GWD [42] and KLD [43] are proposed to use the Gaussian Wasserstein distance and KL divergence to optimize the localization of the bounding boxes, respectively. In this paper, we also build a two-stream oriented object detectors incorporating our proposed $\textup{C}^{2}$ormer to perform extensive experiments.
[…]
@@ III-A Inter-modality Cross-Attention (ICA)
Cross-Correlation Alignment. To find the calibration position between RGB and IR features, we compute the attention map of each feature with respect to another modality features for each modality. The position of maximum weight is the desired aligned features’ position. Thus, we construct the similarity matrices $\tilde{M}\in\mathbb{R}^{HW\times HW}$ of each modality. Specifically, we apply matrix multiplication between query (e.g. $q_{rgb}$) and key (e.g. $k_{ir}$) for different modality descriptors. The RGB and IR matrices can be formulated as follows:
@@
Figure 5: The framework of the $\textup{C}^{2}$ormer-based detectors The outputs of $\textup{C}^{2}$ormer are added into opposite ResNet-50 backbone networks by using an addition operation. For clear illustration, we do not show the FPN structure in this framework.
@@ III-B Adaptive Feature Sampling (AFS)
After that, we generate grids of reference points $p_{rgb}\in\mathbb{R}^{H_{s}\times W_{s}\times 2}$ and $p_{ir}\in\mathbb{R}^{H_{s}\times W_{s}\times 2}$ for the RGB and IR modalities, where $H_{s}=H/s$ and $W_{x}=W/s$ are downsampled from the input featrues’ size by the stride factor $s$. Then, we normalize these points $p_{rgb}$ and $p_{ir}$ to the range $[-1,+1]$, where $(-1,-1)$ represents the top-left position and $(+1,+1)$ represents the bottom-right position. Finally, the sampled features from two modalities ($x_{rgb}$ and $x_{ir}$) are selected at the predicted locations ($p_{rgb}+\Delta p$ and $p_{ir}$) by using the bilinear interpolation sampling function $\varphi(\cdot;\cdot)$ as:
@@
III-C ${C}^{2}$ormer-based Detectors
To evaluate our proposed $\textup{C}^{2}$ormer, we construct a framework of RGB-IR detectors incorporating $\textup{C}^{2}$ormer as shown in Figure 5. Since $\textup{C}^{2}$ormer performs in the feature domain, it can be injected into the single-stage and two-stage object detectors. Specifically, we build two detectors based on $\textup{S}^{2}$-Net [39] and Cascade R-CNN [52] for oriented and horizontal detecion. In this framework, ResNet-50 [53] is utilized as baseline backbone and convolutional layers are used to reduce the channels of RGB-IR features. The input of each $\textup{C}^{2}$ormer is the output of the corresponding stages of the two modalities. Then complementary features output by $\textup{C}^{2}$ormer is added into opposite backbone networks to enhance the representation of modality features. To guarantee the information integrity of the different modality features, we finally feed the sum features of the two modalities into the detection head.
@@ IV-B Implementation Details
For DroneVehicle dataset, we use single-stage oriented object detector—$\rm S^{2}$-Net [39], while for KAIST dataset, we utilize a two-stage horizontal object detector—Cascade R-CNN [52]. We select ResNet50 with FPN (ResNet50-FPN) as the backbone network unless otherwise specified. All detectors are trained for 24 epochs on an NVIDIA TITAN V GPU and optimized using stochastic gradient descent (SGD) algorithm with 0.9 momentum and 0.0001 weight decay. During training, we set two images per batch and initial learning rate is 0.001. At epochs 16 and 22, the rate of learning is divided by 10. We perform augmentation to images, such as flip, crop, and mosaic. During inference, duplicated bounding boxes are removed by conducting the non-maximum suppression (NMS) with 0.3 IoU threshold. All experiments are conducted based on the modified code library MMRotate [56] and MMDetection [57].
Figure 6: Ablation study on different downsampled strides $s$.
@@
Figure 7: Visualization of the intermediate results of the ICA module. The first row shows the example of the red and green points corresponding to the attention map in ${M}_{rgb}$ and ${M}_{ir}$, respectively. The second row visualizes the complementary features (${y}_{ir}$) calculated from the RGB features.

### v3

@@ Abstract
Object detection on visible (RGB) and infrared (IR) images, as an emerging solution to facilitate robust detection for around-the-clock applications, has received extensive attention in recent years. With the help of IR images, object detectors have been more reliable and robust in practical applications by using RGB-IR combined information. However, existing methods still suffer from modality miscalibration and fusion imprecision problems. Since transformer has the powerful capability to model the pairwise correlations between different features, in this paper, we propose a novel Calibrated and Complementary Transformer called $\textbf{C}^{2}\textbf{Former}$ to address these two problems simultaneously. In $\textbf{C}^{2}\textbf{Former}$, we design an Inter-modality Cross-Attention (ICA) module to obtain the calibrated and complementary features by learning the cross-attention relationship between the RGB and IR modality. To reduce the computational cost caused by computing the global attention in ICA, an Adaptive Feature Sampling (AFS) module is introduced to decrease the dimension of feature maps. Because $\textbf{C}^{2}\textbf{Former}$ performs in the feature domain, it can be embedded into existed RGB-IR object detectors via the backbone network. Thus, one single-stage and one two-stage object detector both incorporating our $\textbf{C}^{2}\textbf{Former}$ are constructed to evaluate its effectiveness and versatility. With extensive experiments on the DroneVehicle and KAIST RGB-IR datasets, we verify that our method can fully utilize the RGB-IR complementary information and achieve robust detection results. The code is available at https://github.com/yuanmaoxun/C2Former.git.
@@
RGB-infrared object detection, modality calibration, complementary fusion, multispectral object detection.
@@ I Introduction
Object detection, as a core technology in the computer vision field, has been used in various practical applications, such as autonomous driving, video surveillance and robotics, etc. With the development of convolutional neural network (CNN), the past few years have witnessed significant progress in object detectors [1, 2, 3, 4, 5, 6, 7, 8]. Since these detectors are mainly designed for visible (RGB) images, they cannot cope with the challenges in bad illumination (nighttime) or diverse weather conditions (rain, dust and fog) [9, 10, 11]. For these reasons, detectors exploiting infrared images have been widely adopted as an additional modality for robust object detection [12, 13, 14]. To achieve better performance in the full-time detection task, a growing number of works [15, 16, 17, 18] have been investigated to obtain rich features of objects by integrating visible (RGB) cues and infrared (IR) spectra.
@@
However, there are still two major problems in RGB-IR object detection [20, 19]. (1) One is modality miscalibration as shown in Figure 1(a). Existing methods [21, 22, 16, 23, 24] usually assume that RGB-IR image pairs are perfectly geometrically aligned, and they directly perform the multimodal fusion methods. Actually, after the image registration algorithm, the paired images are just weakly aligned [25, 26]. This is because the RGB-IR image pairs are captured by different sensors with different field-of-views (FoVs) and imaging time, the positions of moving objects may be out of synchronization. (2) Another problem is fusion imprecision as shown in Figure 1(b). In IR modality, the thermal radiation of the vehicle can be important hints for the detector, but none of the RGB images has such cues in poor illumination. The RGB and IR modality features are diverse in terms of object color, texture, and properties, as they are captured in different spectral bands, which brings a big challenge to the fusion strategy. In fact, these two problems are tightly coupled, the misaligned features will directly lead to the imprecision fusion. Thus, these two problems need to be bound together and solved simultaneously.
@@
Recently, some approaches have been introduced to solve these problems in RGB-T detection task. Zhang et al. [25] and Yuan et al. [26] solve the modality miscalibration problem by predicting the proposal offset between two modalities in the RoI head. Although these instance-level prediction solves the object misalignment problem to some extent, they cannot guarantee that the features inside two modality proposals are perfectly aligned. Besides, some cross-modality fusion strategies [10, 27, 15] are also conducted to address the fusion imprecision problem through simple feature concatenation or illumination-aware fusion. However, the inherent complementary information between RGB-IR modality features is still not exploited. More importantly, the above methods do not consider correlating the problems of modality miscalibration and fusion imprecision, but mainly focus on one of them or solve them separately, which will affect the detection performance.
@@
In this paper, referring to the self-attention mechanism in the transformer [28], we formulate the aforementioned two problems into transformer framework and propose a novel Calibrated and Complementary TransFormer-based fusion method called $\textup{C}^{2}$ormer to these two problems simultaneously. Specifically, for each feature point in one modality (Query vector), we first compute the similarity score with all the feature points in the other modality (Key vectors), and then obtain the attention values via the weighted sum versus the features in the other modality (Value vectors). These attention values have two properties: (1) they incorporate the global context information from the other modality, therefore, can be regarded as a better complementary feature than the single local feature point to the original modality. Thus, we can simply concatenate these attention values with the original feature point to conduct an effective feature fusion. (2) The most relevant feature points in the other modality contribute the most to the attention values, which implies that the calibration is automatically achieved thanks to the similarity score. In this way, we simultaneously perform the calibration and fusion operations in a unified framework. Based on this idea, we design an Inter-modality Cross-Attention (ICA) module. An illustration result of the ICA module is given in Figure 2.
@@
Directly computing attention across global feature maps has high computational costs. For that, it needs a feature sampling to reduce the dimension of feature maps. However, because of the miscalibration, an uniform sampling strategy for both modalities may omit meaningful features. To tackle this issue, we present an Adaptive Feature Sampling (AFS) strategy. In detail, we predict an offset versus the foreground object across modalities through a deviation network, and then generate the corresponding sampling strategy according to this offset for different modalities. In this way, we can ensure the integrity of important regions in the sampled feature maps. In actual usage, the given feature maps are first down-sampled by AFS module, and then processed by ICA module to obtain the complementary features, which are finally concatenated with the given feature maps to obtain the fusion features.
@@
The remainder of the article is organized as follows. In the next section, we review the scientific literature related to our proposed approach. Section III analyzes the problem and describes our approach. Experimental results are presented in Section IV, along with the comparison with state-of-the-art methods, and in Section V we conclude with a discussion of our contribution.
@@ II-A Multi-modal Transformers
Since the transformer model [28] has shown its powerful ability to capture long-range dependencies, multi-modal transformers [29, 30, 31, 32, 33, 34, 35] have also been explored. To fuse different modality features, CMSA [36] feeds the words and images into the transformer to construct a joint multimodal feature representation for referring image segmentation and BrT [32] uses a transformer to allow interactive learning by adding a point-to-patch projection between images and point clouds. Similarly, Liu et al. [29] and Xiao et al. [33] design a triplet transformer and feature exchange transformer respectively to detect the salient objects in RGB-D images. However, these methods perform feature fusion by directly concatenating or summing the multi-modalities to feed the transformer. They do not consider the feature misalignment problem, which will lead to a limited representation ability of the fused features.
@@
Different from above approaches, Pixel-BERT [37] considers the modality misalignment problem and aligns vision and language semantics at the pixel and text level using self-supervised learning. Besides, Wang et al. [38] leverage the hierarchical alignment to explore the detailed and diverse characteristics in text-video retrieval task. Recently, Wang et al. [39] aggregate RGB and depth cues through projecting alignment features from other modalities at the token level. Even though these methods consider the feature misalignment problem, the modality complementarity is not further exploited in the fusion process.
@@
Above approaches attempt to utilize the transformer to address multi-modal problems. However, compared with these multi-modal tasks (words-images, texts-videos and point clouds-images), our $\textup{C}^{2}$ormer is designed specific to the RGB-T object detection task and we are the first to apply transformer to this task. Since the problems of modality miscalibration and fusion imprecision are unique and intractable in RGB-T image pairs, we are inspired by the core idea of the self-attention mechanism and formulate these two problems into transformer to establish their correlation. Thus, these two problems are combined and solved simultaneously, which is of great significance in RGB-T object detection task.
II-B Aerial Object Detection
@@
Aerial object detection refers to using rotated bounding box to represent building detectors, which can provide more accurate location of objects. Xia et al. [40] construct a large-scale object detection benchmark with oriented annotations, named DOTA. Existing methods of building detectors based on horizontal bounding boxes are plagued by multiple objects of interest in one anchor. Therefore, many researches about aerial object detection [41, 42, 43, 44, 45, 46] have been presented to alleviate this problem. Ma et al. [41] construct R-RPN to detect aerial objects using rotated proposals. Jiang et al. present $\rm{R^{2}}$ [42], which has a horizontal region of interest (RoI) that is leveraged to predict both horizontal and rotated boxes. To avoid a large number of anchors, RoI Transformer [43] is proposed to transform the horizontal RoI in $\rm{R^{2}}$ into a rotated one. To speed up the inference time of detectors, some works [47, 48, 49, 50] explore one-stage and anchor-free oriented object detectors. $\rm R^{3}$et [47] and $\rm S^{2}$-Net [48] adopt alignments between horizontal receptive fields and rotated anchors to get a better feature representation. Recently, GWD [51] and KLD [52] are proposed to use the Gaussian Wasserstein distance and KL divergence to optimize the localization of the bounding boxes, respectively.
In recent years, utilizing RGB and infrared images for aerial object detection has also received widespread attention. Sun et al. [53] provide a benchmark dataset called DroneVehicle for RGB-IR vehicle detection. On the basis of this dataset, a novel and lightweight detector named CMAFF [54] has been proposed to perform multispectral feature fusion. To further improve vehicle detection performance at nighttime, Gao et al. [55] present a vehicle detection-driven RGB and infrared image fusion method called GF-detection. Furthermore, Yuan et al. [26] observe the weak misalignment problem in DroneVehicle and introduce a TSRA module to calibrate two modality features. Recently, the auxiliary super-resolution branch proposed by [56] has been introduced into RGB-IR objecrt detection, which can effectively distinguish small objects. In this paper, we also build the two-stream aerial object detector incorporating our proposed $\textup{C}^{2}$ormer to perform complementary feature fusion.
@@ III-A Inter-modality Cross-Attention (ICA)
Cross-Correlation Alignment. To calibrate the features between RGB and IR modalities, we compute attention maps between the modality features relative to each other. we first perform the inner product operation between two modality features through matrix multiplication. Then, we normalize the result to obtain the similarity scores. After that, the most relevant feature points (aligned feature) in the other modality have the highest scores. Through above process, we construct the similarity matrices $\tilde{M}\in\mathbb{R}^{HW\times HW}$ of each modality. Specifically, we apply matrix multiplication between query (e.g. $q_{rgb}$) and key (e.g. $k_{ir}$) for different modality descriptors. The RGB and IR matrices can be formulated as follows:
@@
Figure 5: The framework of the $\textup{C}^{2}$ormer-based detectors. Our $\textup{C}^{2}$ormer consists of ICA and AFS, where the input features are first reduced in feature dimension through the pre-module AFS and then output the aligned and fused features by the ICA module. The outputs of $\textup{C}^{2}$ormer are added into opposite ResNet-50 backbone networks by using an addition operation. For clear illustration, we do not show the FPN structure in this framework. Same as the last layer operation, we add the features output by each stage of the two modalities to obtain multi-scale features and finally input them into the FPN.
@@ III-B Adaptive Feature Sampling (AFS)
After that, we generate grids of reference points $p_{rgb}\in\mathbb{R}^{H_{s}\times W_{s}\times 2}$ and $p_{ir}\in\mathbb{R}^{H_{s}\times W_{s}\times 2}$ for the RGB and IR modalities, where $H_{s}=H/s$ and $W_{s}=W/s$ are downsampled from the input featrues’ size by the stride factor $s$. Then, we normalize these points $p_{rgb}$ and $p_{ir}$ to the range $[-1,+1]$, where $(-1,-1)$ represents the top-left position and $(+1,+1)$ represents the bottom-right position. Finally, the sampled features from two modalities ($x_{rgb}$ and $x_{ir}$) are selected at the predicted locations ($p_{rgb}+\Delta p$ and $p_{ir}$) by using the bilinear interpolation sampling function $\varphi(\cdot;\cdot)$ as:
@@
III-C ${C}^{2}$ormer-based Object Detectors
To evaluate our proposed $\textup{C}^{2}$ormer, we design a framework of RGB-IR detectors incorporating $\textup{C}^{2}$ormer as shown in Figure 5. Since $\textup{C}^{2}$ormer performs in the backbone network, different detection heads can be used after it. Respectively, we construct one-stage and two-stage detectors based on $\textup{S}^{2}$-Net [48] and Cascade R-CNN [65] for oriented and horizontal detection respectively. In this framework, ResNet-50 [66] is utilized as baseline backbone and convolutional layers are used to reduce the channels of RGB-IR features. The input of each $\textup{C}^{2}$ormer is the output of the corresponding stages of the two modalities. Then complementary features output by $\textup{C}^{2}$ormer is added into opposite backbone networks to enhance the representation of modality features. To guarantee the information integrity of the different modality features, we finally feed the sum features of the two modalities into the object detection head.
@@ III-C ${C}^{2}$ormer-based Object Detectors
Figure 6: Ablation study on different downsampled strides $s$.
@@ IV-B Implementation Details
For DroneVehicle dataset, we use single-stage oriented object detector—$\rm S^{2}$-Net [48], the size of initial anchors is set to [4, 8, 16] with a single ratio 1.0, while for KAIST dataset, we utilize a two-stage horizontal object detector—Cascade R-CNN [65], the size of initial anchors is [4, 8, 16, 32] with a single ratio 2.43. For RPN settings in Cascade R-CNN, we select top-2000 proposals from each image to guarantee the recall value. We select ResNet50 with FPN (ResNet50-FPN) as the backbone network unless otherwise specified. All detectors are trained for 24 epochs on an NVIDIA TITAN V GPU and optimized using stochastic gradient descent (SGD) algorithm with 0.9 momentum and 0.0001 weight decay. During training, we set two images of size $640\times 512$ per batch and initial learning rate is 0.001. At epochs 16 and 22, the rate of learning is divided by 10. We perform augmentation to images, such as flip, crop, and mosaic. During inference, both detectors are set up the same as in the training phase and the duplicated bounding boxes are removed by conducting the non-maximum suppression (NMS) with 0.3 IoU threshold. All experiments are conducted based on the modified code library MMRotate [68] and MMDetection [69].
@@
Figure 7: Visualization of the intermediate results of the ICA module. The first row shows the example of the red and green points corresponding to the attention map in ${M}_{rgb}$ and ${M}_{ir}$, respectively. The second row visualizes the complementary features ${y}_{ir}$ and ${y}_{rgb}$ calculated from the RGB and IR features respectively.